# ConsistencyDet: Robust Object Detector with Denoising Paradigm of Consistency Model

### v1

@@ Abstract
Object detection, a quintessential task in the realm of perceptual computing, can be tackled using a generative methodology. In the present study, we introduce a novel framework designed to articulate object detection as a denoising diffusion process, which operates on perturbed bounding boxes of annotated entities. This framework, termed ConsistencyDet, leverages an innovative denoising concept known as the Consistency Model. The hallmark of this model is its self-consistency feature, which empowers the model to map distorted information from any temporal stage back to its pristine state, thereby realizing a “one-step denoising” mechanism. Such an attribute markedly elevates the operational efficiency of the model, setting it apart from the conventional Diffusion Model. Throughout the training phase, ConsistencyDet initiates the diffusion sequence with noise-infused boxes derived from the ground-truth annotations and conditions the model to perform the denoising task. Subsequently, in the inference stage, the model employs a denoising sampling strategy that commences with bounding boxes randomly sampled from a normal distribution. Through iterative refinement, the model transforms an assortment of arbitrarily generated boxes into the definitive detections. Comprehensive evaluations employing standard benchmarks, such as MS-COCO and LVIS, corroborate that ConsistencyDet surpasses other leading-edge detectors in performance metrics. Our code is available at https://github.com/qingjiuchina/ConsistencyDet.
@@
Object Detection, Consistency Model, Denoising Paradigm, Self-consistency, Box-renewal
1 Introduction
Object detection stands as a cornerstone task in the realm of computer vision, tasked with the prediction of both positional data and categorical identities for objects within each image [1, 2, 3]. It serves as the foundation for a wide array of applications, encompassing instance segmentation, pose estimation, action recognition, object tracking, and the detection of visual relationships[4, 5, 6, 7, 8]. Initial object detection methods centered on strategies employing sliding windows and region proposals to delineate object candidate regions, utilizing surrogate techniques for regression and classification [9, 10]. While these methods demonstrated some level of effectiveness, they were impeded by the manual design of candidate selection processes, which hampered their adaptability in complex visual contexts. The introduction of anchor boxes offered increased flexibility, enabling models to better accommodate objects of diverse scales and aspect ratios through pre-established anchor points [11, 12]. Despite this improvement, reliance on a priori information persisted, limiting the universality of these methods across various datasets and scenes. In recent years, the field of object detection has shifted from reliance on empirical object priors towards the utilization of learnable object queries or proposals [13, 14]. A seminal development in this direction was the Detection Transformer (DETR), which marked a paradigm shift by employing learnable object queries, thereby recasting object detection as an end-to-end learning challenge [13]. This innovative strategy eradicated the need for hand-crafted components, garnering significant interest in the research community. Concurrently, the Transformer architecture has offered novel perspectives, registering impressive achievements across a spectrum of visual recognition tasks. This movement fostered the emergence of query-based detection models [15, 16], introducing a wave of ingenuity and breadth to the field of object detection. Notably, Transformer models have excelled beyond conventional CNN-based detectors, particularly distinguishing themselves in the domain of small object detection (SOD) techniques.
@@ 1 Introduction
Figure 1: Comparisons of denoising strategies of Diffusion Model and Consistency Model for object detection. Object detection can be regarded as a denoising diffusion process from noisy boxes to object boxes. In Diffusion Model, $q(\cdot|\cdot)$ is the diffusion process and $p_{\theta}(\cdot|\cdot)$ is the reverse process with stepwise denoising operation. In Consistency Model, $f_{\theta}(\cdot,\cdot)$ represents one-step denoising process.
Recently, Diffusion Model [17, 18], also known as a score-based generative model, has demonstrated remarkable effectiveness in a variety of domains, including image generation [19], image segmentation [20], and object detection [21]. A defining characteristic of the Diffusion Model is its iterative sampling mechanism, which methodically reduces noise from an initially random vector. This approach can be adeptly applied to object detection by starting with a set of stochastic bounding boxes. The process begins with entirely random boxes, free from learnable parameters that require training optimization. Through iterative refinement, the method aims to incrementally improve the accuracy of the bounding boxes’ positions and scales. The ultimate goal is to achieve exhaustive delineation of objects within focused categories. This ”noise-to-box” methodology eliminates the need for heuristic object priors or trainable queries, thereby simplifying the selection of object candidates and fostering progress in the development of the object detection.
Building on the principles of the Diffusion Model, DiffusionDet has outperformed various existing detectors, even surpassing the paradigm established by the Transformer model [21]. Yet, its stepwise denoising process, mirroring the iterations of noise addition, imposes constraints on both flexibility and computational efficiency. To render the model viable for real-world applications, these constant iterations require further optimization. Addressing this issue, we propose an innovative approach with the Consistency Model [22], herein referred to as ConsistencyDet. Our model is deeply informed by the foundational concepts of DiffusionDet. A comparative analysis of the denoising strategies employed by the Consistency Model and the Diffusion Model is illustrated in Fig. 1. Distinctly, the self-consistency property of the Consistency Model enables a one-step denoising process, significantly enhancing execution efficiency. Consequently, the number of denoising iterations can be dramatically reduced while maintaining the integrity of detection accuracy.
@@
Figure 2: Consistency Model undergoes training process to establish a mapping, that brings points along any trajectory of the PF ODE back to the origin of that trajectory [22].
As depicted in Fig. 2, our method builds upon the ordinary differential equation (ODE) framework for probability flow (PF) as utilized in the continuous-time model of DiffusionDet [23]. These models effectively direct sample paths, enabling a smooth transition from the initial data distribution to a tractable noise distribution. ConsistencyDet is distinctive in that it maps any given point from an arbitrary time step back to the origin of its trajectory. Thanks to the self-consistency feature of the model, points along the same trajectory correspond to the same initial point. This innovative approach allows for the generation of data samples, originated by the initial points of ODE trajectories, by transforming random noise vectors through a single evaluation of the network. It is worth noting, however, that the Box-renewal operation, which is conducted by amalgamating new proposals with the remaining outputs from ConsistencyDet across various time steps, does improve sample quality, but this comes at an additional computational expense. Nevertheless, the proposed ConsistencyDet achieves cost-effective iterative sampling, akin to the process employed by DiffusionDet.
The proposed ConsistencyDet model innovatively incorporates Gaussian noise [24] into the center coordinates and sizes of actual bounding boxes within images, thereby generating noisy boxes. These noisy boxes are then employed to extract features from the regions of interest (RoI) [10], which are delineated on the comprehensive output feature maps produced by the backbone encoder, including ResNet-50 [25], ResNet-101 [25], and Swin Transformer [26]. Subsequently, the RoI features are inputted into the detection decoder, which aims to accurately predict the ground truth (GT) boxes devoid of noise. As a result, ConsistencyDet demonstrates the capability to infer actual bounding boxes from randomized boxes, effectively fulfilling the object detection task. During the inference phase, ConsistencyDet generates bounding boxes by inversely applying the learned diffusion process.
Moreover, the performance of the proposed ConsistencyDet model has been rigorously assessed on the challenging datasets of MS-COCO and LVIS v1.0 [27, 28]. Supported by a variety of backbone architectures, including ResNet-50 [25], ResNet-101 [25], and Swin-Base [26], ConsistencyDet exhibits commendable performance. Our work contributes to the field in the following ways:
  • We conceptualize object detection as a generative denoising process and propose a novel methodological approach. In contrast to the established paradigm in DiffusionDet, which employs an equal number of iterations for noise addition and removal, our method represents a substantial advancement in enhancing the efficiency of the detection task.
  • In the proposed ConsistencyDet, we have engineered a noise addition and removal paradigm that does not impose specific architectural constraints, thereby allowing for flexible parameterization with a variety of neural network structures. This design choice significantly augments the model’s practicality and adaptability for diverse applications.
  • In crafting the loss function for the proposed ConsistencyDet, we aggregate the individual loss values at time steps $t$ and $t-1$ subsequent to the model’s predictions to compute the total loss. This methodology guarantees that the mapping of any pair of adjacent points along the temporal dimension to the axis origin maintains the highest degree of consistency. This attribute mirrors the inherent self-consistency principle central to the Consistency Model.
The remainder of this paper is structured as follows: Section 2 offers a concise review of the evolution of object detection models, examines the extensive applications of Diffusion Models, and discusses the foundational principles of the Consistency Model. Subsequently, Section 3 delineates the specific methodologies employed for noise addition and removal within the Consistency Model, elucidates the model’s architecture, and provides essential details pertaining to the training and sampling methodologies. Section 4 details the empirical findings obtained from evaluating ConsistencyDet and conducts a comparative analysis against other leading models in the field. The paper concludes with Section 5, which encapsulates the salient features of the newly proposed ConsistencyDet and contemplates avenues for prospective research.
2 Related works and nomenclature
2.1 Object detection
@@
Object detection has undergone significant evolution across various phases. In the initial stages, researchers employed traditional image processing techniques, endeavoring to detect objects by employing methods such as edge detection and feature extraction. These methods, however, encountered considerable challenges when dealing with complex scenes and variations in illumination. With the advent of deep learning, seminal frameworks such as R-CNN and Fast R-CNN [9] have enhanced detection capabilities in terms of speed and accuracy. This was achieved through the integration of innovative concepts like Region Proposal Networks and RoI Pooling. Subsequent developments led to the emergence of single-stage detectors, including YOLO and SSD, which offered improvements in real-time performance, facilitating rapid end-to-end object detection. More recently, the introduction of attention-based Transformer methods, exemplified by DETR [13], has propelled object detection to unprecedented levels of performance, enabling more precise object localization by leveraging global visual context modeling. Although these detection methods have demonstrated noteworthy effectiveness, there remains considerable potential for further advancement. In the present work, we introduce an innovative detection approach that iteratively refines the position and size of bounding boxes. This refinement process employs a series of noisy iterations, culminating in the bounding boxes precisely encompassing the target object.
2.2 Diffusion Model for perception tasks
@@
The Diffusion Model [24, 23], a subset of deep generative models, has emerged as a powerful tool that originates from random distribution samples and gradually reconstructs the expected data through a denoising process. This model has recently achieved remarkable success in a range of fields, including computer vision [18], natural language processing [29], audio signal processing [30], and interdisciplinary applications [31], as highlighted in recent surveys [32]. In the context of object detection, the Diffusion Model has been adapted into a detector termed DiffusionDet [21], which reframes the object detection challenge as a set prediction problem. This involves the assignment of object candidates to GT boxes. DiffusionDet represents a groundbreaking application of the Diffusion Model to the field of object detection. Building upon the foundations of DiffusionDet, our work seeks to optimize the balance between detection accuracy and computational speed. We aim to enhance detection efficiency through a single-step processing approach, while preserving the essential benefits derived from iterative sampling
{threeparttable}
@@
$T$
Number of total time steps
$t$
Current time step
$t_{r}$
A random time step in the range $[0,T]$
$\Delta t$
Time step interval for sampling
[…]
$x/y$-axis coordinate of the $i$-th box’s center point
[…]
Width / Height of the $i$-th box
[…]
$\alpha_{t}/\sigma_{t}$
Parameter in Denoiser at the $t$-th time step
$\theta$
Model parameter
$F_{\theta}(\cdot,\cdot)$
A designed free-form deep neural network
$\mathcal{N}(\cdot,\cdot)$
Normal distribution
$f_{\theta}(\cdot,\cdot)$
Final answer for Consistency Model
$c_{skip/out/in}(\cdot)$
Calculation factor for $f_{\theta}$
$d(\cdot,\cdot)$
Distance function
$\eta$
Learning rate
$\mu$
EMA decay rate
$\Phi(\cdot,\cdot;\phi)$
ODE solver
$\lambda(\cdot)$
A positive weighting function
$\mathcal{L}$
Total loss function in training process
$\mathcal{L}_{cls/L1/giou}$
Focal / L1 / GIoU loss item
$\lambda_{cls/L1/giou}$
Weight for Focal / L1 / GIoU loss item
$\sigma_{max/min}$
Maximum / Minimum threshold of noise parameter
$\sigma_{data}$
Noise parameter between $\sigma_{min}$ and $\sigma_{max}$
$\epsilon$
Randomly generated Gaussian noise
$\rho$
Scale factor of generated noise
$B_{t}$
Random noise at the $t$-th time step in sampling
$r(\cdot)$
Generate random noise with given dimensions
$E(\cdot)$
Image feature extraction with backbone network
$P_{c}(\cdot,\cdot)$
Prediction of Consistency Model in each time step
$nms(\cdot,\cdot)$
Non-max suppression (NMS) operation
[…]
Threshold of NMS operation
[…]
Threshold of Box-renewal operation
$dcm(\cdot,\cdot)$
Decoder of ConsistencyDet with head network
$ddm(\cdot,\cdot)$
Denoiser of DiffusionDet with head network
$conc(\cdot,\cdot)$
Concatenate function
$n_{ss}$
Number of sampling steps
$n_{p}$
Number of total proposed boxes
$n_{r}$
Number of current proposed boxes
$x_{s}$
Padded box information at time axis origin
$x_{t}$
Noised box information at the $t$-th time step
$x_{b}$
Predicted box information in each time step
$x_{0}$
Predicted box information at time axis origin
$x_{box/cls}$
Predicted object’s box coordinate / category
AP
Average Precision
[…]
Average Precision at 50% / 75% IoU
[…]
Average Precision for small / median / large objects
[…]
Average Precision for rare / common / frequent categories
@@
2.3 Consistency Model
The Diffusion Model is predicated on an iterative generation process, which tends to result in sluggish execution efficiency, thus curtailing its applicability in real-time scenarios. To surmount this limitation, OpenAI has unveiled the Consistency Model, an innovative category of generative models capable of rapidly producing high-quality samples without necessitating adversarial training regimes. The Consistency Model facilitates swift one-step generation, yet it retains the option for multi-step sampling as a means to navigate the trade-off between computational efficiency and the caliber of generated samples. Furthermore, it introduces the capability for zero-shot [33] data manipulation, encompassing tasks such as image restoration, colorization, and super-resolution, obviating the need for task-specific training.
The Consistency Model can be cultivated through a distillation pre-training regimen derived from existing Diffusion Models or alternatively as a standalone generative model. This work formally acknowledges this capability and, for the inaugural time, incorporates the Consistency Model within the domain of object detection, hereby designated as ConsistencyDet.
2.4 Nomenclature

### v5

# ConsistencyDet: A Few-step Denoising Framework for Object Detection Using the Consistency Model

@@ Abstract
Object detection, a quintessential task in the realm of perceptual computing, can be tackled using a generative methodology. In the present study, we introduce a novel framework designed to articulate object detection as a denoising diffusion process, which operates on the perturbed bounding boxes of annotated entities. This framework, termed ConsistencyDet, leverages an innovative denoising concept known as the Consistency Model. The hallmark of this model is its self-consistency feature, which empowers the model to map distorted information from any time step back to its pristine state, thereby realizing a “few-step denoising” mechanism. Such an attribute markedly elevates the operational efficiency of the model, setting it apart from the conventional Diffusion Model. Throughout the training phase, ConsistencyDet initiates the diffusion sequence with noise-infused boxes derived from the ground-truth annotations and conditions the model to perform the denoising task. Subsequently, in the inference stage, the model employs a denoising sampling strategy that commences with bounding boxes randomly sampled from a normal distribution. Through iterative refinement, the model transforms an assortment of arbitrarily generated boxes into definitive detections. Comprehensive evaluations employing standard benchmarks, such as MS-COCO and LVIS, corroborate that ConsistencyDet surpasses other leading-edge detectors in performance metrics. Our code is available at https://anonymous.4open.science/r/ConsistencyDet-37D5.
@@
Object Detection, Consistency Model, Denoising Paradigm, Self-consistency, Box-renewal.
I Introduction
Object detection is a fundamental task in computer vision that predicts both positional data and categorical identities for objects in images [1, 2, 3]. It supports various applications, including instance segmentation, pose estimation, and action recognition [4, 5, 6, 7, 8]. Early methods relied on sliding windows and region proposals for defining candidate regions, using surrogate techniques for regression and classification [9, 10]. While effective, these approaches were limited by manual candidate selection, affecting adaptability in complex contexts. The introduction of anchor boxes enhanced flexibility for diverse object scales and aspect ratios [11, 12], although reliance on prior information restricted their applicability across datasets. Recently, the field has shifted towards learnable object queries or proposals [13, 14], with significant advancements such as the Detection Transformer (DETR), which employs learnable queries to recast detection as an end-to-end learning problem [13]. This approach eliminated the need for hand-crafted components and led to impressive results in various visual recognition tasks, especially in small object detection (SOD).
@@ I Introduction
Figure 1: Comparisons of denoising strategies of the Diffusion Model and Consistency Model for object detection. Object detection can be regarded as a denoising diffusion process from noisy boxes to object boxes. In the Diffusion Model, $q(\cdot|\cdot)$ is the diffusion process and $p_{\theta}(\cdot|\cdot)$ is the reverse process with a stepwise denoising operation. In the Consistency Model, $f_{\theta}(\cdot,\cdot)$ represents a one-step denoising process.
Diffusion Model [15, 16], a score-based generative model, has shown remarkable effectiveness in image generation [17], segmentation [18], and object detection [19]. Its iterative sampling mechanism systematically reduces noise from an initially random vector. In object detection, this process begins with stochastic bounding boxes, refining accuracy through iterations to delineate objects effectively. This “noise-to-box” methodology obviates the need for heuristic priors, simplifying object candidate selection. Based on Diffusion Model principles, DiffusionDet has outperformed existing detectors, even surpassing the Transformer model paradigm [19].
However, its incremental denoising process limits flexibility and computational efficiency, requiring further optimization for practical applications. We propose an innovative framework, ConsistencyDet, building on the foundational principles of DiffusionDet [20]. As shown in Fig. 1, a comparative analysis reveals that the self-consistency property of the Consistency Model enables few-step denoising, significantly improving computational efficiency while reducing the required iteration count without sacrificing detection accuracy. In Fig. 2, our approach leverages the ordinary differential equation (ODE) framework for probability flow (PF) from DiffusionDet [21], facilitating a smooth transition from data distributions to noise distributions. ConsistencyDet uniquely maps points from arbitrary time steps back to their trajectory origin, enabling efficient sample generation by transforming random noise vectors into structured outputs with minimal network evaluations.
@@
Figure 2: Consistency Model undergoes training process to establish a mapping that brings points along any trajectory of the PF ODE back to the origin of that trajectory [20].
Furthermore, ConsistencyDet effectively integrates Gaussian noise [22] into the coordinates and sizes of bounding boxes, creating noisy boxes for feature extraction from regions of interest (RoI) [10]. These RoI features, which are obtained from output maps produced by deep backbone networks like ResNet-50 [23], are then fed into a detection decoder to predict noise-free ground truth boxes. Consequently, ConsistencyDet can accurately infer the true bounding boxes from randomly generated ones, generating them through an inverse diffusion process with minimal computational cost and high efficiency.
The performance of ConsistencyDet has been rigorously evaluated on challenging datasets like MS-COCO and LVIS v1.0 [24, 25], demonstrating commendable results across various backbones, including ResNet-50, ResNet-101, and Swin-Base [23, 26]. Our contributions are as follows:
  • Based on the self-consistency characteristic of the Consistency Model, in the design of the loss function, we accumulate the loss values at time steps $t$ and $t-1$ to compute the overall loss. This ensures that the mapping results of adjacent points along the temporal axis to the origin remain stable, thereby enhancing the model’s “few-step denoising” capability.
  • Our work introduces the principles of the Consistency Model to the task of object detection for the first time, proposing a novel detection method and establishing a well-designed pipeline for it, termed ConsistencyDet. Our work offers a new perspective and research direction for the field of object detection.
  • Based on the unique characteristic of self-consistency, ConsistencyDet enables the model to efficiently map distorted information from any temporal stage back to its original state, thereby achieving a highly effective “few-step denoising” mechanism. Compared to other object detection methods based on the diffusion model, this distinctive feature significantly enhances the model’s overall inference speed and computational efficiency.
II Related works and nomenclature
II-A Object detection
@@
Object detection has undergone significant evolution across various phases. In the initial stages, researchers employed traditional image processing techniques, endeavoring to detect objects by employing methods such as edge detection and feature extraction. These methods, however, encountered considerable challenges when dealing with complex scenes and variations in illumination. With the advent of deep learning, seminal frameworks such as R-CNN and Fast R-CNN [9] have significantly enhanced detection capabilities in terms of speed and accuracy. This was achieved through the integration of innovative concepts like Region Proposal Networks and RoI Pooling. Subsequent developments led to the emergence of single-stage detectors, including YOLO and SSD, which offered considerable improvements in real-time performance, facilitating rapid end-to-end object detection. More recently, the introduction of attention-based Transformer methods, exemplified by DETR [13], has propelled object detection to unprecedented levels of performance, enabling more precise object localization by leveraging global visual context modeling. Although these detection methods have demonstrated noteworthy effectiveness, there remains considerable potential for further advancement. In the present work, we introduce an innovative detection approach that iteratively refines the position and size of bounding boxes. This iterative refinement process employs a series of noisy iterations, culminating in the bounding boxes precisely encompassing the target object.
II-B Diffusion Model for perception tasks
@@
The Diffusion Model [27, 22, 28, 21] is a powerful and highly effective generative model that reconstructs data from random noise through a denoising process and has shown impressive success in various fields like computer vision [16] and natural language processing [29]. In object detection, it has been adapted into DiffusionDet [19], which reinterprets detection as a set prediction task by assigning object candidates to ground truth (GT) boxes. Building upon DiffusionDet, our work aims to further enhance detection efficiency by introducing a novel single-step method that preserves the benefits of iterative sampling. This approach optimizes the balance between detection accuracy and computational speed, providing a streamlined yet highly effective and computationally efficient solution for complex object detection tasks.
@@
$T$ Number of total time steps
$t$ Current time step
$t_{r}$ A random time step in the range $[0,T]$
$\Delta t$ Time step interval for sampling
$\textit{c}_{x}^{\textit{i}}/\textit{c}_{y}^{\textit{i}}$ $x/y$-axis coordinate of the $i$-th box’s center point
$\textit{w}^{\textit{i}}/\textit{h}^{\textit{i}}$ Width / Height of the $i$-th box
$\textit{b}^{\textit{i}}$ $(\textit{c}_{x}^{\textit{i}},\textit{c}_{y}^{\textit{i}},\textit{w}^{\textit{i}},\textit{h}^{\textit{i}})$ of the $i$-th box
$\alpha_{t}/\sigma_{t}$ Parameter in Denoiser at the $t$-th time step
$\theta$ Model parameter
$F_{\theta}(\cdot,\cdot)$ A designed free-form deep neural network
$\mathcal{N}(\cdot,\cdot)$ Normal distribution
$f_{\theta}(\cdot,\cdot)$ Final answer for Consistency Model
$c_{skip/out/in}(\cdot)$ Calculation factor for $f_{\theta}$
$d(\cdot,\cdot)$ Distance function
$\eta$ Learning rate
$\mu$ EMA decay rate
$\Phi(\cdot,\cdot;\phi)$ ODE solver
$\lambda(\cdot)$ A positive weighting function
$\mathcal{L}$ Total loss function in training phase
$\mathcal{L}_{cls/L1/giou}$ Focal / L1 / GIoU loss item
$\lambda_{cls/L1/giou}$ Weight for Focal / L1 / GIoU loss item
$\sigma_{max/min}$ Maximum / Minimum threshold of noise parameter
$\sigma_{data}$ Noise parameter between $\sigma_{min}$ and $\sigma_{max}$
$\epsilon$ Randomly generated Gaussian noise
$\rho$ Scale factor of generated noise
$B_{t}$ Random noise at the $t$-th time step in sampling
$r(\cdot)$ Generate random noise with given dimensions
$E(\cdot)$ Image feature extraction with backbone network
$P_{c}(\cdot,\cdot)$ Prediction of Consistency Model in each time step
$nms(\cdot,\cdot)$ Non-max suppression (NMS) operation
$N_{th}$ Threshold of NMS operation
$B_{th}$ Threshold of Box-renewal operation
$dcm(\cdot,\cdot)$ Decoder of ConsistencyDet with head network
$ddm(\cdot,\cdot)$ Denoiser of DiffusionDet with head network
$conc(\cdot,\cdot)$ Concatenate function
$n_{ss}$ Number of sampling steps
$n_{tr}$ Number of total proposed boxes in training phase
$n_{p}$ Number of total proposed boxes in inference
$n_{r}$ Number of current proposed boxes
$x_{s}$ Padded box information at time axis origin
$x_{t}$ Noised box information at the $t$-th time step
$x_{b}$ Predicted box information in each time step
$x_{0}$ Predicted box information at time axis origin
$x_{box/cls}$ Predicted object’s box coordinate / category
AP Average Precision
AP50/75 Average Precision at 50% / 75% IoU
APs/m/l Average Precision for small / median / large objects
APr/c/f Average Precision for rare / common / frequent categories
@@
II-C Consistency Model
The Diffusion Model is predicated on an iterative generation process, which tends to result in sluggish execution efficiency, thus curtailing its applicability in real-time scenarios. To address this limitation, OpenAI has unveiled the Consistency Model, an innovative category of generative models capable of rapidly producing high-quality samples without necessitating adversarial training regimes. The Consistency Model facilitates swift one-step generation, yet it retains the option for multi-step sampling as a means to navigate the trade-off between computational efficiency and the caliber of generated samples. Furthermore, it introduces the capability for zero-shot data manipulation [30], encompassing tasks such as image restoration, colorization, and super-resolution, obviating the need for task-specific training. The Consistency Model can be cultivated through a distillation pre-training regimen derived from existing Diffusion Models or alternatively as a standalone generative model. This work formally acknowledges this capability and, for the inaugural time, incorporates the Consistency Model within the domain of object detection, hereby designated as ConsistencyDet.
II-D Nomenclature